# User Location Disclosure Fails to Deter Overseas Criticism but Amplifies Regional Divisions on Chinese Social Media

### v1

@@ Abstract
We examine the behavioral impact of a user location disclosure policy implemented on Sina Weibo, China’s largest microblogging platform, using a high-frequency, real-time dataset of uncensored user engagement with 165 leading government and media accounts. Leveraging a natural experiment result from the platform’s sudden rollout of location tagging on April 28, 2022, we compare millions of time-stamped observations of user behavior in the comment sections of these accounts before and after the policy change. Although the policy appeared intended to deter overseas users from spreading information deemed harmful by the regime, we find no reduction in their engagement. Instead, the policy sharply reduced domestic users’ willingness to comment on posts about local issues outside their own provinces. This effect was especially pronounced among out-of-province commenters and disproportionately curtailed criticisms. Using large language models, we further show that location disclosure triggered a rise in regionally discriminatory replies, which in turn heightened the perceived risk of cross-provincial engagement and reshaped the norms of online participation. Our findings suggest that authoritarian regimes can reinforce censorship not only through top-down control, but by mobilizing social cleavages—here, regional divisions—to suppress dissent and fragment public discourse.
@@ 1 Introduction
In October 2021, the Cyberspace Administration of China released draft regulations that sparked public outcry and drew intense international media attention (Reuters, 2022; The New York Times, 2022). These regulations required online service providers to “prominently display IP address locations of user accounts,” specifying that platforms must show the province for domestic users and the country or region for overseas users. While the Chinese government justified the policy as a measure to curb online misinformation and counter foreign influence, it drew criticism for further restricting freedom of speech for Chinese internet users. On March 17, 2022, less than a month after Russia began its invasion of Ukraine, Sina Weibo, China’s largest microblogging platform, revealed users’ IP locations in accordance with the new regulations. Weibo added a location tag to each post or comment related to the Russo-Ukrainian war, claiming it aimed to minimize misinformation and ensure content authenticity. However, many believe the real intent was to restrict overseas users from sharing information that contradicted the government’s stance. [^1] For example, Weibo’s announcement highlights a user’s IP location in Ukraine, subtly insinuating foreign users as potential purveyors of misinformation. In a later press conference, A speaker from the Taiwan Affairs Office of the State Council even made the purpose of the policy clear: “to pinpoint and neutralize individuals from Taiwan who could be intentionally stirring the pot in cross-strait relations.” See http://www.gwytb.gov.cn/m/speech/202205/t20220511_12435159.htm. Accessed on May 2024. At noon on April 28, Weibo abruptly disclosed each user’s location on every post and comment. Other major Chinese social media companies, including Douyin, Toutiao, Zhihu, Kuaishou, and Xiaohongshu, soon followed suit with similar policies. By the end of 2022, all social media platforms in China had adopted the user’s IP disclosure policy.
@@
This paper examines the causal impact of the location disclosure policy implemented on Weibo. Authoritarian regimes use a range of tools to manage online discourse, including censorship, surveillance, and increasingly, user information disclosure. Prior work shows that Chinese authorities remove content linked to collective action while tolerating some criticism (King et al., 2013a; Lorentzen, 2014; Roberts, 2018). These cyberspace regulations deter participation through both targeted repression and ambient monitoring (Esarey and Xiao, 2011; Pan et al., 2023). In addition, social media platforms like Weibo function as tools for state messaging, information gathering, and controlled public engagement (Chen and Xu, 2017a, b; Qin et al., 2017). With platform cooperation, the state has adopted user-level labeling tools—such as real-name registration and location disclosure—to steer user behavior without direct censorship. These policies make identity attributes visible to others and can change user behavior. Despite their growing adoption, empirical evidence on the behavioral effects of such policies remains limited, largely due to data and identification challenges. A smaller body of research suggests that visible identity cues—like regional tags—can heighten social divisions and incivility (Peng, 2021; Chen et al., 2021; Yu and Margolin, 2024; Guo et al., 2023). However, little is known about their causal impact on user behavior and online discourse.
Studying the effects of user information disclosure presents several empirical challenges. First, data accuracy is a central concern. In heavily censored environments like Weibo, post-censorship data can bias inference—changes in critical comments may reflect shifts in censorship intensity rather than actual behavior. Prior work shows that around 3% of sensitive posts are removed within 30 minutes and 90% within 24 hours (King et al., 2013b), while non-sensitive posts remain online, creating survivorship bias. Public datasets, such as Weiboscope and Weibo-Cov (Weiboscope Project, ; Nguyen, 2020), often amplify this issue: many rely on crawling schemes that oversample high-follower accounts or are assembled post-censorship using narrow keyword lists. Moreover, basic metadata, such as user location, often rely on users’ self-designated tags, which may introduce systematic geographic error (Qin et al., 2018). These limitations make it difficult to construct samples that reliably capture authentic user behavior.
Second, measurement is often limited. Most studies using existing Weibo datasets retain only original posts and high-level reposts (e.g., Lu et al., 2021; Qin et al., 2024; Meng et al., 2025), omitting detailed engagement data such as the volume and content of engagement, and nested threads—where much of the deliberation, conflict, and behavioral response to government policies takes place. Without such data, it is difficult to analyze how conversations evolve within comment sections. Moreover, researchers often lack information about which users are subject to information disclosure and what specific details are revealed, making it difficult to identify who is treated.
Third, causal identification remains challenging. To our knowledge, no experimental studies exist, in part due to ethical constraints. Observational studies often rely on platform-generated signals such as deletion flags or user tags, but these are governed by opaque criteria and applied unevenly. Without knowing how the platform selects users for disclosure, unobserved confounding is a serious concern. Most datasets rely on low-frequency scraping and do not track accounts or posts over time, which limits their usefulness for studying behavioral dynamics over time.
We address these empirical challenges through a real-time, high-frequency data collection strategy that continuously monitors a set of prominent government-affiliated Weibo accounts and their comment sections. This approach yields a highly granular, longitudinal dataset at both the account and post levels. By scraping data roughly every five minutes, we capture most comments before they are censored or removed, minimizing the influence of post-publication moderation. In addition to post-level engagement metrics, we track top comments and their replies, which are ranked by the platform primarily based on likes and responses. This allows us to observe how discussions evolve and how the top comment section updates over time. The dataset spans a sudden, unannounced policy change, enabling credible before-and-after comparisons using an interrupted time series design. Its longitudinal structure supports within-account comparisons across the policy threshold, while repeated scraping of the first-page comment section captures user discourse over an extended period.
@@ 2 Data and Measurement
We collaborated with industry experts to design a real-time data collection system that monitored 165 prominent Weibo accounts affiliated with the Chinese government. These include official accounts of government offices and state media outlets—such as newspapers, television stations, and major media outlets—identified in a series of reports on institutional influence in 2021.222These reports include multiple editions of Government Weibo Influence Reports (政务微博影响力报告), Sina Government Weibo Reports (新浪政务微博报告), and Weibo Trending Topics Data Reports (微博热搜榜数据报告). See Table LABEL:table:account_list in the SM for a full list of accounts we collected data from. We focused on these accounts because, following the government’s 2013 crackdown on key opinion leaders (Pan et al., 2025), the comment sections of these accounts—due to high traffic and visibility—have become important spaces for public expression and policy discussion. These accounts reach large audiences and maintain high engagement, making them well-suited for observing shifts in public sentiment. Our aim is to capture both the content posted by these accounts and the real-time public engagement their posts generate.
@@ 2.1 High-frequency Weibo Data
We conducted real-time monitoring of each account’s timeline (i.e., their posts), each post’s first-page comment section (displaying up to the top 20 comments), and associated engagement metrics at both the post and comment levels. We focus on the first-page comments because they were instantly available to readers when a post is clicked and draw the most user attention and interaction. As shown in the top-left panel of Figure 1(a), our system scanned all monitored accounts, as well as recorded posts and comments, every five minutes. This allowed us to collect longitudinal data at both the account and post levels. When a new post appeared, the system initiated tracking and began recording data on the post and its top comments. Snapshots were captured every five minutes for the first 24 hours, and once every day for the following ten days. While most engagement occurred within the first few hours (see Figure S4 in the SM), the extended tracking window enabled us to capture delayed or atypical activity. Figure 1(b) illustrates the variables collected through scraping, including each post’s content, publication time, and engagement metrics (number of reposts, likes, and comments), as well as the first-page of the comment section (top 20 comments). For each comment and their replies, we also recorded the associated IP location tag. Figure 1(c) and Figure 1(d) show the numbers of new posts, and new comments collected within a 48-hour window, respectively. Figures 1(e) and 1(f) show that, on average, comments appearing on and after the second page receive minimal engagement.
@@ 2.2 Timeline and Location Tags
The location tags reflect users’ IP addresses at the time of commenting, although we do not observe or record the IPs themselves. Overseas users are labeled by country or region, while domestic users are tagged at the provincial level (e.g., Shanghai, Jiangsu, Shaanxi). [^3] A potential concern is that users may mask their IP addresses using VPNs. In the SM, we show that fewer than 1% of users changed location tags during the study period. VPN usage among domestic users is rare and typically serves to switch between domestic and foreign IPs, not between provinces. To our knowledge, commercially available VPN services in China do not support intra-national IP masking, as all provinces operate within the same national firewall. These patterns suggest that provincial-level location tags are generally reliable. Excluding users whose location tags have changed does not alter on our findings.
@@
Post content: positive, negative, or neutral event Post LLM-assisted annotation + human validation $0.721$
@@
Comment stance: critical of the post or not Comment Supervised BERT classifier + human validation $0.783$
Comment receiving regionally discriminatory replies Comment LLM-assisted annotation + human validation $0.835$
@@ 2.3 Key variables and Measurement
We also classify post types based on their content. International posts are those that mention one of 123 countries or regions outside China. We construct the list of 123 countries and regions (see Table S4 in the SM) by extracting all foreign country and region names that appeared in Weibo’s post-treatment IP location tags. Local posts reference one of China’s 31 provinces. We focus on provincial references, rather than prefectures or counties, because IP location tags are reported at the country level for overseas users and the provincial level for domestic users. Posts are further categorized as depicting a positive, negative, or neutral event (from the perspective of the Chinese government) using a large language model (LLM), followed by human validation.
@@
To detect regional discrimination in replies, we use LLM-based annotation, validated by human coders, to identify whether a reply contains discriminatory content based on the commenter’s location. We also track the frequency with which a user’s IP location is mentioned in replies as an indirect measure of the salience of geographic identity in the discourse. Table 1 summarizes these variables and their corresponding measurement strategies. Table S3 in the SM provides the summary statistics.
@@ 2.4 Protecting Privacy
All data used in this study are publicly available. Nevertheless, we take additional steps to protect the privacy of commenters, many of whom are likely ordinary citizens in China. First, we do not collect any personally identifiable information, such as commenters’ IP addresses. The location tags we use are limited to the provincial level and are too coarse to identify individuals. Second, we mask all user identifiers—including usernames, avatars, and URLs—by replacing them with randomly assigned IDs to ensure complete anonymity. Third, all data are stored on a secure server at the corresponding author’s institution, with access restricted to authorized personnel only (see Table S2 for the protocol.)
@@ 3 Identification Strategy
Because an interrupted time series design is analogous to a regression discontinuity design with time as the forcing variable (Hausman and Rapson, 2018), we estimate $\tau_{\text{SAITT}}$ using local linear regression on either side of $c$, clustering standard errors at the account level. Bandwidths are selected using both the MSE-optimal procedure in rdrobust (Calonico et al., 2014) and cross-validation. We show in the SM (Figures S12 to S17) that our results are robust to a wide range of bandwidth choices.
@@ 4.1 Revealing IP locations failed to deter overseas users but reduced domestic engagement
The policy was designed to deter participation by overseas users with “malicious intent,” who were believed to interfere in domestic discourse by spreading misinformation and expressing views that challenge the official narrative. The assumption was that once users’ IP locations became public, domestic users would be better equipped to judge the credibility of online content, thereby reducing the effectiveness of malicious foreign propaganda. At the same time, overseas users might reduce their engagement out of concern about being discredited or attacked as part of a “malicious foreign force.” To evaluate whether the policy has achieved its goal, we first distinguish between posts related to international affairs and those that are not. Posts by government and media accounts on international affairs are defined as those that explicitly mention at least one foreign location from the aforementioned country list; all other posts are classified as non-international, which include both posts covering domestic issues and general content without a clear reference to a foreign country.
@@
One alternative explanation is that domestic users, masking their locations via Virtual Private Networks (VPNs), contributed to this surge. Yet this is unlikely: we find that IP location tags remained unchanged for 99.1% of users before and after the policy was implemented, indicating minimal switching. Furthermore, analysis in Figure S25 in the SM shows that overseas users continued to engage with sensitive topics such as the Russian invasion of Ukraine, despite facing increased attacks from domestic commenters, likely provoked by the newly visible location tags following the policy implementation.
Moreover, we find no decline in comments on non-international posts (red line in Figure 2(a)) by international users, further weakening the deterrence hypothesis. The red line remains consistently lower than the blue line, but has much narrower confidence intervals. This is because overseas users are more likely to comment on posts related to international affairs; however, such posts are less frequently made by propaganda accounts compared to non-international posts, which explains the wider confidence intervals around the blue line.
@@
In Figure 2(c), we further examine the change in comment volume by domestic users on posts related to non-international issues, where a sharp decline followed the introduction of the policy. We divide these posts into two categories: those concerning local issues (purple lines)—defined as posts that mention 31 province names in mainland China (see Table S5 in the SM for the list)—and those addressing other, non-local topics (red lines). We define local issues in this manner because the IP location tags were assigned at the provincial level. The results show that the overall decline in the volume of comments on non-international content in Figure 2(b) is primarily driven by a drop in engagement with posts about local issues. In fact, the average number of comments per post on local issues by domestic users dropped sharply from 8.21 to approximately 0.42 after the policy was implemented. [^4] Consistent with this finding, Figure S27 in the SM shows that aggregate engagement metrics, including the total number of comments, likes, and reports, on posts about local issues also declined following the policy. However, using the aggregate metrics, we cannot distinguish whether this decrease is driven by domestic or international users. In contrast, the volume of comments on non-local issues (posts without mentioning a province name) remains relatively stable before and after the policy’s implementation.
@@
To assess this possibility, we examine whether there was any change in the content of the posts themselves. If, for instance, government or media accounts began posting fewer negative or politically sensitive stories after the policy was introduced, that might explain the decline in comments. However, Figure 2(d) shows that the proportion of local posts covering positive, neutral, and negative events remained stable before and after the policy’s implementation. These negative events include major accidents, scandals, and COVID-lockdown. In fact, we show in Figure S6 in the SM that comments declined across posts featuring both positive and negative events. This suggests that the drop in engagement was not driven by changes in what was being posted, but by changes in who chose to respond.

### v2

@@ Abstract
We examine the behavioral effects of a user location disclosure policy implemented by Sina Weibo, China’s largest microblogging platform, using a high-frequency dataset of uncensored user engagement—including tens of thousands of comments—on 165 prominent government and media accounts. Exploiting the platform’s abrupt rollout of IP-based location tags on April 28, 2022, we compare user behavior in comment sections before and after the policy change. Although the policy was publicly justified as a measure to curb misinformation and counter foreign influence, we find no decline in participation by overseas users. Instead, it significantly reduced domestic engagement with local issues outside users’ home provinces, particularly among critical comments. Evidence suggests this effect was not driven by generalized fear or concerns about credibility, but by a rise in regionally discriminatory replies that increased the social cost of cross-provincial engagement. Our findings indicate that identity disclosure tools can produce unintended consequences by activating existing social divisions in ways that reinforce state control without direct censorship.
@@ 1 Introduction
In October 2021, the Cyberspace Administration of China released draft regulations requiring platforms to display users’ IP-based locations—by province for domestic users and by country for overseas users. The Chinese government framed the policy as a measure to curb misinformation and foreign interference, but critics viewed it as another step to constrain speech online (Reuters, 2022; The New York Times, 2022). On March 17, 2022, less than a month after the Russian invasion of Ukraine, Sina Weibo, China’s largest microblogging site, began tagging the locations of users posting about the Russo-Ukrainian war, claiming it aimed to minimize misinformation and ensure content authenticity. However, many believe the real intent was to restrict overseas users from sharing information that contradicted the government’s stance (see Supporting Information, SI). At noon on April 28, Weibo abruptly expanded the policy to all posts and comments. Other major platforms quickly followed. By the end of 2022, location disclosure had become a standard feature of Chinese social media.
@@
This paper examines the causal impact of the location disclosure policy implemented on Weibo. Authoritarian regimes use a range of tools to manage online discourse, including censorship (King et al., 2013a; Lorentzen, 2014; Roberts, 2018), surveillance (Esarey and Xiao, 2011; Pan et al., 2023), and orchestrated engagement (Chen and Xu, 2017a, b; Qin et al., 2017). Increasingly, they also rely on platform-enabled identity disclosure—such as real-name registration or location tagging—to steer behavior without direct repression. These tools make identity attributes visible to others and may influence behavior through reputational or social pressures. While prior research suggests that regional cues can heighten incivility and deepen social divisions (Peng, 2021; Chen et al., 2021; Yu and Margolin, 2024; Guo et al., 2023), their causal impact on online participation remains poorly understood. This study helps fill that gap. Beyond the authoritarian context, this study contributes to a broader literature on how platform architecture shapes user behavior. Prior research shows that design features—such as source trustworthiness ratings (Papakyriakopoulos et al., 2022), state-affiliation tags (Liang et al., 2022; Nassetta and Gross, 2020; Bradshaw et al., 2024), and feed-ranking algorithms (Flaxman et al., 2016; Guess et al., 2023)—can influence exposure, sharing, and engagement, even without changing users’ beliefs. Our findings add to this work by showing that simple identity cues, like location tags, can also suppress participation and segment discourse, highlighting the broader behavioral consequences of platform design choices.
Study the effects of user information disclosure faces empirical challenges. First, data accuracy is difficult to ensure in hihgly censored environments like Weibo, where behavioral signals may reflect censorship rather than genuine user response. Prior work shows that 3% of sensitive posts are deleted within 30 minutes and 90% within 24 hours (King et al., 2013b), introducing survivorship bias. Public datasets such as Weiboscope and Weibo-Cov (Weiboscope Project, 2018; Nguyen, 2020) often rely on post-censorship crawling, which overrepresents high-follower accounts, or uses narrow keyword lists. Location data, when available, typically rely on self-reported tags subject to systematic misclassification (Qin et al., 2018). Second, measurement is limited: most studies preserve only original posts and high-level reposts (e.g., Lu et al., 2021; Qin et al., 2024; Meng et al., 2025), omitting comment threads where user reactions to policy often emerge. Information on which users are subject to disclosure—and what is revealed—is typically missing, making it difficult to measure treatment exposure. Third, credible causal inference is rare. Experimental studies are constrained by ethical concerns, while observational approaches rely on platform-generated signals, such as deletion flags or user tags; these signals are usually governed by opaque and inconsistently applied rules, leading to potential confounding. Moreover, most datasets are scraped at low frequency and do not track accounts or posts over time, which limits their value for studying behavioral dynamics.
To overcome these challenges, we implemented a high-frequency data collection strategy that continuously monitored a set of prominent government-affiliated Weibo accounts and their comment sections. Scraping at approximately five-minute intervals allowed us to capture most comments before censorship or removal, reducing survivorship bias and enabling us to track evolving discussions with high temporal resolution. We focused on top comments and their replies—ranked by the platform based on engagement—to observe how public discourse unfolds in visible spaces. The dataset spans the sudden, unannounced rollout of the location disclosure policy on Weibo, providing a natural interruption for before-and-after comparisons. Its panel structure supports within-account comparisons, while repeated snapshots of the comment section capture shifts in user engagement over time.
@@ 2 Data and Measurement
We collaborated with industry experts to design a real-time data collection system that monitored 165 prominent Weibo accounts affiliated with the Chinese government. These include official accounts of government offices and state media outlets—such as newspapers, television stations, and major media outlets—identified in a series of reports on institutional influence in 2021.111These reports include multiple editions of Government Weibo Influence Reports (政务微博影响力报告), Sina Government Weibo Reports (新浪政务微博报告), and Weibo Trending Topics Data Reports (微博热搜榜数据报告). See Table LABEL:table:account_list in the SI for a full list of accounts we collected data from. We focused on these accounts because, following the government’s 2013 crackdown on key opinion leaders (Pan et al., 2025), the comment sections of these accounts—due to high traffic and visibility—have become important spaces for public expression and policy discussion. These accounts reach large audiences and maintain high engagement, making them well-suited for observing shifts in public sentiment. Our aim is to capture both the content posted by these accounts and the real-time public engagement their posts generate.
@@ 2.1 High-frequency Weibo Data
We conducted real-time monitoring of each account’s timeline (i.e., their posts), each post’s first-page comment section (displaying up to the top 20 comments), and associated engagement metrics at both the post and comment levels. We focus on the first-page comments because they were instantly available to readers when a post is clicked and draw the most user attention and interaction. As shown in the top-left panel of Figure 1(a), our system scanned all monitored accounts, as well as recorded posts and comments, every five minutes. This allowed us to collect longitudinal data at both the account and post levels. When a new post appeared, the system initiated tracking and began recording data on the post and its top comments. Snapshots were captured every five minutes for the first 24 hours, and once every day for the following ten days. While most engagement occurred within the first few hours (see Figure S4 in the SI), the extended tracking window enabled us to capture delayed or atypical activity. Figure 1(b) illustrates the variables collected through scraping, including each post’s content, publication time, and engagement metrics (number of reposts, likes, and comments), as well as the first-page of the comment section (top 20 comments). For each comment and their replies, we also recorded the associated IP location tag. Figure 1(c) and Figure 1(d) show the numbers of new posts, and new comments collected within a 48-hour window, respectively. Figures 1(e) and 1(f) show that, on average, comments appearing on and after the second page receive minimal engagement.
@@ 2.2 Timeline and Location Tags
The location tags reflect users’ IP addresses at the time of commenting, although we do not observe or record the IPs themselves. Overseas users are labeled by country or region, while domestic users are tagged at the provincial level (e.g., Shanghai, Jiangsu, Shaanxi). [^2] A potential concern is that users may mask their IP addresses using VPNs. in the SI, we show that fewer than 1% of users changed location tags during the study period. VPN usage among domestic users is rare and typically serves to switch between domestic and foreign IPs, not between provinces. To our knowledge, commercially available VPN services in China do not support intra-national IP masking, as all provinces operate within the same national firewall. These patterns suggest that provincial-level location tags are generally reliable. Excluding users whose location tags have changed does not alter on our findings.
@@
Post content: positive, negative, or neutral event Post LLM-assisted annotation + human validation 0.7210.7210.721
@@
Comment stance: critical of the post or not Comment Supervised BERT classifier + human validation 0.7830.7830.783
Comment receiving regionally discriminatory replies Comment LLM-assisted annotation + human validation 0.8350.8350.835
@@ 2.3 Key variables and Measurement
We also classify post types based on their content. International posts are those that mention one of 123 countries or regions outside China. We construct the list of 123 countries and regions (see Table S4 in the SI) by extracting all foreign country and region names that appeared in Weibo’s post-treatment IP location tags. Local posts reference one of China’s 31 provinces. We focus on provincial references, rather than prefectures or counties, because IP location tags are reported at the country level for overseas users and the provincial level for domestic users. Posts are further categorized as depicting a positive, negative, or neutral event (from the perspective of the Chinese government) using a large language model (LLM), followed by human validation.
@@
To detect regional discrimination in replies, we use LLM-based annotation, validated by human coders, to identify whether a reply contains discriminatory content based on the commenter’s location. We also track the frequency with which a user’s IP location is mentioned in replies as an indirect measure of the salience of geographic identity in the discourse. Table 1 summarizes these variables and their corresponding measurement strategies. Table S3 in the SI provides the summary statistics.
@@ 2.4 Protecting Privacy
All data used in this study are publicly available. Nevertheless, we take additional steps to protect the privacy of commenters, many of whom are likely ordinary citizens in China. First, we do not collect any personally identifiable information, such as commenters’ IP addresses. The location tags we use are limited to the provincial level and are too coarse to identify individuals. Second, we mask all user identifiers—including usernames, avatars, and URLs—by replacing them with randomly assigned IDs to ensure complete anonymity. Third, all data are stored on a secure server at the corresponding author’s institution, with access restricted to authorized personnel only (see Table S2 in the SI for details.)
@@ 3 Identification Strategy
Because an interrupted time series design is analogous to a regression discontinuity design with time as the forcing variable (Hausman and Rapson, 2018), we estimate $\tau_{\text{SAITT}}$ using local linear regression on either side of $c$, clustering standard errors at the account level. Bandwidths are selected using both the MSE-optimal procedure in rdrobust (Calonico et al., 2014) and cross-validation. We show in the SI (Figures S12 to S17) that our results are robust to a wide range of bandwidth choices.
@@ 4.1 Revealing IP locations failed to deter overseas users but reduced domestic engagement
The policy was designed to deter participation by overseas users with “malicious intent,” who were believed to interfere in domestic discourse by spreading misinformation and expressing views that challenge the official narrative. The assumption was that once users’ IP locations became public, domestic users would be better equipped to judge the credibility of online content, thereby reducing the effectiveness of malicious foreign propaganda. At the same time, overseas users might reduce their engagement out of concern about being discredited or attacked as part of a “malicious foreign force” (see SI). To evaluate whether the policy has achieved its goal, we first distinguish between posts related to international affairs and those that are not. Posts by government and media accounts on international affairs are defined as those that explicitly mention at least one foreign location from the aforementioned country list; all other posts are classified as non-international, which include both posts covering domestic issues and general content without a clear reference to a foreign country.
@@
One alternative explanation is that domestic users, masking their locations via Virtual Private Networks (VPNs), contributed to this surge. Yet this is unlikely: we find that IP location tags remained unchanged for 99.1% of users before and after the policy was implemented, indicating minimal switching. Furthermore, analysis in Figure S25 in the SI shows that overseas users continued to engage with sensitive topics such as the Russian invasion of Ukraine, despite facing increased attacks from domestic commenters, likely provoked by the newly visible location tags following the policy implementation.
Moreover, we find no decline in comments on non-international posts (red line in Figure 2(a)) by international users, which further weakens the deterrence hypothesis. The red line remains consistently lower than the blue line, but has much narrower confidence intervals. This is because overseas users are more likely to comment on posts related to international affairs; however, such posts are less frequently made by propaganda accounts compared to non-international posts, which explains the wider confidence intervals around the blue line.
@@
In Figure 2(c), we further examine the change in comment volume by domestic users on posts related to non-international issues, where a sharp decline followed the introduction of the policy. We divide these posts into two categories: those concerning local issues (purple lines)—defined as posts that mention 31 province names in mainland China (see Table S5 in the SI for the list)—and those addressing other, non-local topics (red lines). We define local issues in this manner because the IP location tags were assigned at the provincial level. The results show that the overall decline in the volume of comments on non-international content in Figure 2(b) is primarily driven by a drop in engagement with posts about local issues. In fact, the average number of comments per post on local issues by domestic users dropped sharply from 8.21 to approximately 0.42 after the policy was implemented. [^3] Consistent with this finding, Figure S27 in the SI shows that aggregate engagement metrics, including the total number of comments, likes, and reports, on posts about local issues also declined following the policy. However, using the aggregate metrics, we cannot distinguish whether this decrease is driven by domestic or international users. In contrast, the volume of comments on non-local issues (posts without mentioning a province name) remains relatively stable before and after the policy’s implementation.
@@
To assess this possibility, we examine whether there was any change in the content of the posts themselves. If, for instance, government or media accounts began posting fewer negative or politically sensitive stories after the policy was introduced, that might explain the decline in comments. However, Figure 2(d) shows that the proportion of local posts covering positive, neutral, and negative events remained stable before and after the policy’s implementation. These negative events include major accidents, scandals, and COVID-lockdown. In fact, we show in Figure S6 in the SI that comments declined across posts featuring both positive and negative events. This suggests that the drop in engagement was not driven by changes in what was being posted, but by changes in who chose to respond.